OpenStudy (jacobbirddog):

whoever answers this correctly gets a fan!!!! Apples sell for $1.90 per pound, and bananas sell for $0.75 per pound. Troy bought some apples and some bananas. Together they weighed 3.8 pounds, and cost $5.84. Let a represent the number of pounds of apples Troy purchases. Which expression represents the number of pounds of bananas he purchases? A. 3.8a B. 3.8 – a C. a + 3.8 D. a – 3.8

OpenStudy (anonymous):

B, because you take the total pounds, which is 3.80 and subtract it by the weight of the apples to find the weight of the bananas.

OpenStudy (jacobbirddog):

ty! @tn.x680

OpenStudy (anonymous):

You're welcome!

OpenStudy (jacobbirddog):

tn i have another one!

OpenStudy (jacobbirddog):

Carrots sell for $2.10 per pound, and crackers sell for $2.90 per pound. Glen bought some carrots and some crackers. The total weight was 2.3 pounds and cost $6.03. Let p represent the number of pounds of carrots. Which equation represents the situation described? A. –0.8p + 6.03 = 6.67 B. 2.1p + 2.9 = 6.03 C. –0.8p + 6.67 = 6.03 D. 2.1p + 6.03 = 2.9

OpenStudy (anonymous):

B, because carrots are $2.10 per pound, and p represents pounds, so 2.1p would give you the price of the carrots. Crackers are 2.90 so you add that to the previous and set it equal to the total price, 6.03. 2.1p + 2.9 = 6.03